# Instance-level Performance Prediction for Long-form Generation Tasks

### v1

@@ A.3 Additional Dataset Information
APIGen (Liu et al., 2025) Function Calling 8955 / 2985 / 2985 CodeBLEU, CodeJudge (functional correctness, inconsistency)[gemini-2.0-flash-lite]
ASQA (Stelmakh et al., 2022) Question Answering 1689 / 563 / 564 LLM-EVAL (accuracy, informativeness) [gemini-2.0-flash-lite], BERTScore
BigCodeBench (Zhuo et al., 2025) Code Generation 1994 / 665 / 665 CodeBLEU, CodeJudge (functional correctness, inconsistency)[gemini-2.0-flash-lite]
BiGGen (Kim et al., 2025) Open-Ended Generation 847 / 283 / 283 LLM-EVAL (accuracy, informativeness) [gemini-2.0-flash-lite]
@@
Eli5 (Fan et al., 2019) Question Answering 2657 / 886 / 886 BERTScore, LLM-EVAL (accuracy, informativeness) [gemini-2.0-flash-lite]
FActScore-Bio (Min et al., 2023) Fact-Checking 1216 / 406 / 406 FActScore [gpt-3.5-turbo]
MedLFQA (Hosseini et al., ) Medical QA 8855 / 2952 / 2952 BERTScore, LLM-EVAL (accuracy, informativeness) [gemini-2.0-flash-lite]
Opus-100 (EN→ES) (Zhang et al., 2020; Tiedemann, 2012) Machine Translation 3579 / 1194 / 1194 BERTScore, GEMBA_SQM_ref [gemini-2.0-flash-lite]
Summeval (Fabbri et al., 2021) Text Summarization 2809 / 937 / 937 BERTScore, G-Eval (Coherence, Consistency, Fluency, Relevance) [gemini-2.0-flash-lite]
TruthfulQA (Lin et al., 2022) Question Answering 1455 / 486 / 486 BERTScore, LLM-EVAL (accuracy, informativeness) [gemini-2.0-flash-lite]
Table 5: Our study and benchmark cover twelve datasets spanning a range of diverse tasks: code generation, fact-checking, function-calling, machine translation (MT), question-answering (QA), summarization, and others. The original dataset sizes is capped at 5,000 instances (through random sampling), with each instances generating three outputs, we collect maximum 15,000 output-score instance per datasets. Train-dev-test is split by 3:1:1. For each dataset, we also specify one or more task-specific, multi-dimensional and partial-credit evaluation metric used to evaluate generation performance.
@@
APIGEN (Liu et al., 2025) CodeBLEU
ASQA (Stelmakh et al., 2022) BERTScore
BigCodeBench (Zhuo et al., 2025) CodeBLEU
@@
Eli5 (Fan et al., 2019) BERTScore
FActscore-Bio (Min et al., 2023) FActScore
MED LF QA (Hosseini et al., ) BERTScore
Opus-100 (EN to ES) (Zhang et al., 2020; Tiedemann, 2012) BERTScore
Summeval (Fabbri et al., 2021) BERTScore
TruthfulQA (Lin et al., 2022) BERTScore
@@ A.4 Complementary Tables and Figures
Dataset Task-Specific Metric Score
APIGen CodeBLEU 0.32
Judge(Func.) 0.77
Judge(Incon.) 0.56
ASQA BERTScore 0.89
LLMEval(A.) 0.87
LLMEval(I.) 0.84
BiGGen Judge 0.67
BigCodeBench CodeBLEU 0.11
Judge(Func.) 0.77
Judge(Incon.) 0.63
Summeval G-Eval(Coh.) 0.88
G-Eval(Con.) 0.95
G-Eval(Flu.) 0.59
G-Eval(Rel.) 0.79
BERTScore 0.90
Eli5 BERTScore 0.85
LLMEval(A.) 0.72
LLMEval(I.) 0.63
FActScore-Bio FActScore 0.43
Med-LFQA BERTScore 0.84
LLMEval(A.) 0.76
LLMEval(I.) 0.59
Opus-100 GEMBA-SQM_ref 0.87
BERTScore 0.85
TruthfulQA BERTScore 0.87
LLMEval(A.) 0.60
LLMEval(I.) 0.58
Table 7: Llama 3.2 11B performance on 10 long-form datasets with various task-specific evaluation metrics. All scores are normalized by their respective scale to [0,1]. Performance prediction then seeks to predict these gold-standard evaluation scores.
@@
CE-Reg: Consistency-based CE-Reg: Verbalized RF-LLMaaJ Dataset Metric DegMat. Eccen. EigVal. LexSim. NumSet. 1S 2S 4-shot 8-shots 16-shot APIGen Judge(Func.) 0.48 0.38 0.38 0.42 0.34 0.19 0.16 0.25 0.27 0.23 Judge(Incon.) 0.55 0.35 0.30 0.35 0.17 0.09 0.13 0.03 0.01 0.08 CodeBLEU 0.65 0.38 0.41 0.47 0.08 0.02 0.11 0.05 0.04 0.01 ASQA LLMEval(A.) 0.47 0.30 0.43 0.52 0.23 0.41 0.35 0.14 0.14 0.11 LLMEval(I.) 0.38 0.27 0.36 0.42 0.16 0.33 0.27 0.02 0.05 0.08 BERTScore 0.51 0.35 0.53 0.58 0.07 0.34 0.16 0.07 0.05 0.10 BiGGen Judge 0.62 0.26 0.52 0.55 0.10 0.21 0.25 0.26 0.21 0.06 BigCodeBench Judge(Func.) 0.46 0.36 0.41 0.48 0.07 0.11 0.15 0.06 0.01 0.02 Judge(Incon.) 0.33 0.27 0.26 0.35 0.04 0.08 0.13 0.09 0.02 0.04 CodeBLEU 0.56 0.47 0.50 0.56 0.09 0.00 0.05 0.07 0.03 -0.07 Eli5 LLMEval(A.) 0.50 0.18 0.48 0.46 0.31 0.16 0.21 0.12 0.09 0.15 LLMEval(I.) 0.36 0.13 0.31 0.30 0.23 0.17 0.22 0.22 0.24 0.18 BERTScore 0.54 0.32 0.53 0.51 0.06 0.01 0.11 -0.02 -0.04 0.02 Fact-Bio FActScore 0.72 0.25 0.74 0.77 0.27 0.68 0.26 0.47 0.50 0.54 MED LF QA LLMEval(A.) 0.48 0.24 0.49 0.50 0.24 0.31 0.24 0.11 0.13 0.13 LLMEval(I.) 0.44 0.30 0.42 0.43 0.07 0.28 0.24 0.24 0.28 0.27 BERTScore 0.55 0.38 0.52 0.54 0.16 0.22 0.36 0.09 0.04 0.05 Opus-100 GEMBA 0.52 0.37 0.48 0.55 0.29 0.24 0.14 0.09 0.11 0.08 BERTScore 0.66 0.50 0.63 0.65 0.23 0.11 0.15 0.06 0.05 -0.02 Summeval G-Eval(Coh.) 0.38 0.17 0.36 0.35 0.09 0.23 0.02 0.00 0.04 0.02 G-Eval(Con.) 0.29 0.10 0.34 0.30 0.12 0.17 0.02 -0.02 -0.04 -0.07 G-Eval(Flu.) 0.08 0.00 0.08 0.08 0.08 0.04 -0.03 -0.04 0.05 -0.03 G-Eval(Rel.) 0.40 0.13 0.38 0.38 0.16 0.23 0.02 0.02 0.03 0.01 BERTScore 0.56 0.24 0.57 0.52 0.08 0.16 0.00 -0.03 0.04 -0.02 TruthfulQA LLMEval(A.) 0.60 0.24 0.54 0.54 0.12 0.18 0.31 0.06 -0.01 0.12 LLMEval(I.) 0.43 0.21 0.35 0.39 0.14 0.20 0.18 0.15 0.06 0.20 BERTScore 0.60 0.13 0.39 0.46 0.08 0.12 0.11 0.03 0.10 -0.02 AVERAGE 0.49 0.27 0.43 0.46 0.15 0.20 0.16 0.10 0.09 0.08
Table 9: PC Benchmark. We compare two baselines: 1) confidence-based regression (CE-Reg) uses 7 CE features (five consistency-based, two verbalized); and 2) reference-free LLM-as-a-Judge (RF-LLMaaJ), evaluated in 4-, 8-, and 16- settings. For PC, values higher the better. Judge(Func.) = CodeJudge Functional Correctness. Judge(Incon.) = CodeJudge Inconsistency. LLMEval(A.) = LLM Eval Accuracy. LLMEval(I.) = LLMEval Informativeness.

### v2

@@ A.3 Additional Dataset Information
APIGen (Liu et al., 2025) Function Calling 8955 / 2985 / 2985 CodeBLEU, CodeJudge (functional correctness, inconsistency)[gemini-2.5-flash]
ASQA (Stelmakh et al., 2022) Question Answering 1689 / 563 / 564 LLM‑EVAL (accuracy, informativeness) [gemini-2.5-flash], BERTScore
BigCodeBench (Zhuo et al., 2025) Code Generation 1994 / 665 / 665 CodeBLEU, CodeJudge (functional correctness, inconsistency)[gemini-2.5-flash]
BiGGen (Kim et al., 2025) Open-Ended Generation 847 / 283 / 283 BiGGen-Judge [gemini-2.5-flash]
@@
Eli5 (Fan et al., 2019) Question Answering 2657 / 886 / 886 BERTScore, LLM‑EVAL (accuracy, informativeness) [gemini-2.5-flash]
Fact-Bio (Min et al., 2023) Fact-Checking 1216 / 406 / 406 FActScore [gpt-3.5-turbo]
Med LF QA (Hosseini et al., ) Medical QA 8855 / 2952 / 2952 BERTScore, LLM‑EVAL (accuracy, informativeness) [gemini-2.5-flash]
Opus-100 (EN→ES) (Zhang et al., 2020; Tiedemann, 2012) Machine Translation 3579 / 1194 / 1194 BERTScore, GEMBA_SQM_ref [gemini-2.5-flash]
Summeval (Fabbri et al., 2021) Text Summarization 2809 / 937 / 937 BERTScore, G-Eval (Coherence, Consistency, Fluency, Relevance) [gemini-2.5-flash]
TruthfulQA (Lin et al., 2022) Question Answering 1455 / 486 / 486 BERTScore, LLM‑EVAL (accuracy, informativeness) [gemini-2.5-flash]
Table 6: Our study and benchmark cover twelve datasets spanning a range of diverse tasks: code generation, fact-checking, function-calling, machine translation (MT), question-answering (QA), summarization, and others. The original dataset sizes is capped at 5,000 instances (through random sampling), with each instances generating three outputs, we collect maximum 15,000 output-score instance per datasets. Train-dev-test is split by 3:1:1. For each dataset, we also specify one or more task-specific, multi-dimensional and partial-credit evaluation metric used to evaluate generation performance. BiGGen-Bench-Results is the direct assessment evaluation setting.
@@
APIGen (Liu et al., 2025) Judge (Funct.)
ASQA (Stelmakh et al., 2022) LLMEval(A.)
BigCodeBench (Zhuo et al., 2025) Judge (Funct.)
@@
Eli5 (Fan et al., 2019) LLMEval(A.)
Fact-Bio (Min et al., 2023) FActScore
MED LF QA (Hosseini et al., ) LLMEval(A.)
Opus-100 (Zhang et al., 2020; Tiedemann, 2012) GEMBA
Summeval (Fabbri et al., 2021) G-Eval(Rel.)
TruthfulQA (Lin et al., 2022) LLMEval(A.)
@@ A.4 Complementary Tables and Figures
CE-Reg: Consistency-based CE-Reg: Verbalized RF-LLMaaJ Dataset Metric DegMat. Eccen. EigVal. LexSim. NumSet. 1S 2S 4-shot 8-shot 16-shot Judge(Func.) 0.39 0.38 0.40 0.39 0.37 0.39 0.39 0.43 0.43 0.43 Judge(Incon.) 0.11 0.12 0.12 0.12 0.12 0.12 0.12 0.23 0.21 0.20 APIGen CodeBLEU 0.05 0.06 0.06 0.06 0.06 0.06 0.06 0.12 0.11 0.10 ASQA LLMEval(A.) 0.25 0.24 0.25 0.25 0.23 0.23 0.24 0.26 0.26 0.25 LLMEval(I.) 0.17 0.17 0.17 0.17 0.17 0.17 0.18 0.20 0.18 0.20 ASQA BERTScore 0.02 0.02 0.02 0.02 0.03 0.02 0.02 0.08 0.05 0.05 BiGGen Judge 0.30 0.31 0.29 0.28 0.30 0.30 0.31 0.35 0.36 0.35 BigCodeBench Judge(Func.) 0.36 0.34 0.36 0.37 0.35 0.35 0.35 0.48 0.49 0.50 Judge(Incon.) 0.09 0.09 0.09 0.09 0.09 0.09 0.09 0.22 0.19 0.17 BigCodeBench CodeBLEU 0.08 0.09 0.08 0.08 0.10 0.10 0.10 0.14 0.13 0.13 LLMEval(A.) 0.26 0.28 0.26 0.26 0.27 0.28 0.28 0.33 0.33 0.32 LLMEval(I.) 0.17 0.18 0.16 0.17 0.18 0.18 0.18 0.21 0.21 0.21 Eli5 BERTScore 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.06 0.07 0.05 Fact-Bio FActScore 0.20 0.31 0.20 0.22 0.31 0.25 0.32 0.33 0.31 0.34 LLMEval(A.) 0.28 0.29 0.28 0.28 0.28 0.28 0.29 0.34 0.34 0.32 LLMEval(I.) 0.18 0.19 0.18 0.18 0.19 0.18 0.18 0.21 0.21 0.21 MED LF QA BERTScore 0.03 0.03 0.03 0.03 0.03 0.03 0.03 0.09 0.09 0.09 Opus-100 GEMBA 0.21 0.22 0.22 0.21 0.22 0.22 0.23 0.26 0.24 0.24 Opus-100 BERTScore 0.06 0.07 0.07 0.06 0.08 0.08 0.08 0.12 0.12 0.12 G-Eval(Coh.) 0.14 0.12 0.13 0.14 0.12 0.12 0.12 0.16 0.16 0.16 G-Eval(Con.) 0.18 0.15 0.17 0.17 0.14 0.15 0.15 0.19 0.18 0.20 G-Eval(Flu.) 0.05 0.05 0.05 0.06 0.05 0.05 0.05 0.07 0.07 0.07 G-Eval(Rel.) 0.20 0.18 0.20 0.21 0.18 0.18 0.18 0.21 0.22 0.22 Summeval BERTScore 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.06 0.06 0.06 LLMEval(A.) 0.31 0.34 0.30 0.30 0.35 0.34 0.34 0.44 0.43 0.43 LLMEval(I.) 0.27 0.28 0.27 0.26 0.29 0.29 0.29 0.35 0.36 0.34 TruthfulQA BERTScore 0.04 0.04 0.04 0.04 0.04 0.04 0.04 0.15 0.12 0.12 AVERAGE 0.16 0.17 0.16 0.16 0.17 0.17 0.17 0.23 0.22 0.22
Table 8: RMSE Benchmark. We compare two baselines: 1) confidence-based regression (CE-Reg) uses 7 CE features (five consistency-based, two verbalized); and 2) reference-free LLM-as-a-Judge (RF-LLMaaJ), evaluated in 4-, 8-, and 16-shot settings. For RMSE, lower is better. Judge(Func.) = CodeJudge Functional Correctness. Judge(Incon.) = CodeJudge Inconsistency. LLMEval(A.) = LLM Eval Accuracy. LLMEval(I.) = LLMEval Informativeness.
CE-Reg: Consistency-based CE-Reg: Verbalized RF-LLMaaJ Dataset Metric DegMat. Eccen. EigVal. LexSim. NumSet. 1S 2S 4-shot 8-shot 16-shot Judge(Func.) 0.40 0.36 0.31 0.36 0.36 0.17 0.20 0.29 0.31 0.33 Judge(Incon.) 0.46 0.31 0.29 0.34 0.26 0.13 0.17 -0.14 -0.14 -0.10 APIGen CodeBLEU 0.59 0.33 0.38 0.41 0.10 -0.01 0.15 0.08 0.05 0.06 ASQA LLMEval(A.) 0.33 0.24 0.35 0.36 0.30 0.33 0.15 0.24 0.22 0.25 LLMEval(I.) 0.39 0.27 0.38 0.41 0.30 0.28 0.15 0.18 0.28 0.14 ASQA BERTScore 0.51 0.34 0.48 0.57 0.10 0.33 0.23 0.04 0.07 0.11 BiGGen Judge 0.44 0.13 0.44 0.50 0.27 0.24 0.21 0.19 0.08 0.16 BigCodeBench Judge(Func.) 0.34 0.32 0.33 0.33 0.05 0.10 0.04 0.07 0.10 0.08 Judge(Incon.) 0.36 0.35 0.35 0.35 0.10 0.06 -0.01 -0.08 -0.08 -0.08 BigCodeBench CodeBLEU 0.58 0.50 0.56 0.60 0.12 0.05 0.02 0.04 0.09 0.05 LLMEval(A.) 0.48 0.19 0.49 0.50 0.27 0.12 0.21 0.13 0.04 0.10 LLMEval(I.) 0.46 0.20 0.50 0.50 0.26 0.10 0.18 0.19 0.14 0.18 Eli5 BERTScore 0.58 0.27 0.57 0.59 -0.01 -0.03 0.08 0.04 0.06 0.07 Fact-Bio FActScore 0.78 0.31 0.78 0.75 0.30 0.65 0.21 0.45 0.56 0.54 LLMEval(A.) 0.42 0.17 0.43 0.45 0.23 0.23 0.13 0.10 0.08 0.13 LLMEval(I.) 0.42 0.16 0.39 0.43 0.13 0.24 0.16 0.20 0.22 0.24 MED LF QA BERTScore 0.55 0.38 0.52 0.54 0.16 0.22 0.36 0.09 0.04 0.05 Opus-100 GEMBA 0.50 0.36 0.44 0.49 0.28 0.33 0.21 0.16 0.21 0.22 Opus-100 BERTScore 0.66 0.50 0.63 0.65 0.23 0.11 0.15 0.06 0.05 -0.02 G-Eval(Coh.) 0.01 0.05 0.03 0.02 0.16 0.05 0.04 0.06 -0.01 -0.02 G-Eval(Con.) 0.01 0.09 0.10 0.05 0.15 0.08 0.02 0.02 0.03 -0.01 G-Eval(Flu.) 0.09 0.01 0.10 0.05 0.03 0.03 0.03 0.00 -0.02 0.02 G-Eval(Rel.) 0.15 0.09 0.17 0.13 0.14 0.18 0.05 0.04 0.01 0.06 Summeval BERTScore 0.56 0.24 0.57 0.52 0.08 0.16 0.00 -0.03 0.04 -0.02 LLMEval(A.) 0.53 0.27 0.55 0.55 0.08 0.25 0.22 0.04 0.08 0.06 LLMEval(I.) 0.48 0.24 0.45 0.49 0.06 0.14 0.15 0.07 0.04 0.12 TruthfulQA BERTScore 0.60 0.13 0.39 0.46 0.08 0.12 0.11 0.03 0.10 -0.02 AVERAGE 0.43 0.25 0.41 0.42 0.17 0.17 0.13 0.09 0.10 0.08
Table 9: Pearson Correlation Benchmark. We compare two baselines: 1) confidence-based regression (CE-Reg) uses 7 CE features (five consistency-based, two verbalized); and 2) reference-free LLM-as-a-Judge (RF-LLMaaJ), evaluated in 4-, 8-, and 16-shot settings. For RMSE, lower is better. Judge(Func.) = CodeJudge Functional Correctness. Judge(Incon.) = CodeJudge Inconsistency. LLMEval(A.) = LLM Eval Accuracy. LLMEval(I.) = LLMEval Informativeness.